OpenStudy (anonymous):

what are the roots of the equation : x^2-10x-21=0 ?

OpenStudy (saifoo.khan):

Quadratic formula?

OpenStudy (anonymous):

complete the square!

OpenStudy (anonymous):

if the coefficient of the middle term is even easiest to complete the square then you have less to do at the end

OpenStudy (anonymous):

\[x^2-10x=21\] \[(x-5)^2=21+5^2=46\] \[x-5=\pm\sqrt{46}\] \[x=5\pm\sqrt{46}\]